OpenStudy (anonymous):

Determine the extreme values:

OpenStudy (anonymous):

\[\huge y= \frac{ e^{x} }{ 1+e^{x} }, x \epsilon [0,4]\]

OpenStudy (fozia):

values maximum, minimum

OpenStudy (anonymous):

yes, how tho?

OpenStudy (anonymous):

i find the derivative first and then what?

OpenStudy (fozia):

first fint derivative that gives the result y'=ex/(1+ex)2 then eqaute it to zero gives ex=0

OpenStudy (anonymous):

\[\huge y'=\frac{ e^x }{ (1+e^x) }\]

OpenStudy (fozia):

take 2nd derivatv now put the value of x =0 in 2nd derivativ if relt is postv then its min otherwise max

OpenStudy (anonymous):

@fozia wait, you and I got different derivatives

OpenStudy (anonymous):

@fozia , you have correctly identified the first derivative, but it will never be equal to zero. Therefore, there is no need to take the second derivative. It is sufficient to just analyze the first derivative, realize that it is always positive, and that the original function is always increasing, so it has no extreme values.

OpenStudy (anonymous):

check if the function is increasing or decreasing?

OpenStudy (anonymous):

Here's a graph to show you that it is always increasing

OpenStudy (anonymous):

how do i check if its inc/dec?

OpenStudy (anonymous):

its first derivative is always a positive, hence it is increasing for all [0.4]. the minimum value will be at x=0, maximu will be at x=4.

OpenStudy (anonymous):

Read my first post. That will tell you. That numerator of the first derivative is always positive. So is the denominator.

OpenStudy (anonymous):

No extreme values for domain of all x. min and max though at the endpoints for the given problem domain.

OpenStudy (anonymous):

That's because "e" is a positive base, so any exponent you put on it will result in a positive number.

OpenStudy (anonymous):

Do you see that fozia's derivative is correct? That's where you should start.

OpenStudy (anonymous):

Use the quotient rule to get the derivative. Do you need help with that?

OpenStudy (anonymous):

@burhan101 I have 2 pending questions posted for you here.

OpenStudy (anonymous):

@tcarroll010 im trying to figure out how that derivative is right

OpenStudy (anonymous):

& yes i am using the quotient rule

OpenStudy (anonymous):

\[\huge \frac{ (1+e^x)(e^x)-(e^x)(e^x)}{ (1+e^x)^2 }\]

OpenStudy (anonymous):

\[\frac{ (1 + e ^{x})e ^{x} - e ^{x}e ^{x} }{ (1 + e ^{x})^{2} } = \frac{ e ^{x} + e ^{2x} - e ^{2x} }{ (1 + e ^{x})^{2} }\]

OpenStudy (anonymous):

In the numerator, you are left with only e^x and the denominator is the original denominator squared.

OpenStudy (anonymous):

ohhhh is see it

OpenStudy (anonymous):

So, there is no need to take the second derivative. It is sufficient to just analyze the first derivative, realize that it is always positive, and that the original function is always increasing, so it has no extreme values, if your domain is all x. But your domain is limited, so you have your min and max at the x-value endpoints.

OpenStudy (anonymous):

but dont i plug in 0 and 4 in f(x) ?

OpenStudy (anonymous):

So, min at: e^0 / (1 + e^0) max at: e^4 / (1 + e^4)

OpenStudy (anonymous):

All good now, @burhan101 ?

OpenStudy (anonymous):

yes thanks !!

OpenStudy (anonymous):

uw! Good luck in all of your studies and thx for the recognition! @burhan101

OpenStudy (anonymous):

thank you very much, you do explain things very clear !

OpenStudy (anonymous):

I just like to help. Kind words on your part!

OpenStudy (anonymous):

For this function \[\huge f(x)= 2\sin4x+3 ; [0, \pi]\] i get the derivative as \[\huge f'(x)=8\cos4x \] To find the extreme values, do i plug in the endpoints into f(x) ? @tcarroll010

OpenStudy (dan815):

YO BURHAANN MY BRO

OpenStudy (dan815):

are u done with this equation

OpenStudy (anonymous):

dannn:) ! no im not, do i plug the endpoints into the function ?

OpenStudy (dan815):

explain the meaning of an extreme value to me

OpenStudy (anonymous):

an absolute mx or an absolute min

OpenStudy (dan815):

ok

OpenStudy (dan815):

so that means the slope there is 0

OpenStudy (anonymous):

yes

OpenStudy (dan815):

but u must check the end points too incase its just a local max or local min there

OpenStudy (dan815):

first solve f'(x) = 0

OpenStudy (dan815):